if I conduct an absorption experiment using packed bed, with constant solvent rate and variable gas rate, will coefficient of mass transfer of liquid phase be identical throughout the experiment, even though gas rate is variable?
Answer #1

I

In the process of the absorption liquid absorbs some of the contaminants(say A) from the gas.
In the given above diagram we can see that Xb is the mole fraction of the A in the bulk liquid.
Xi is the mole fraction of the A in the liquid phase near the interface.
Yi is the mole fraction of the gas in the gaseous phase.
Xb is the mole fraction of the A in the bulk gas phase.

We can assume the process to be interface absorption.

Flux of A from the liquid to the gas is given as = Kx(Xb-Xi)
Since,the amount of A from the liquid to the interface will be equal to that of the amount from the interface to the gas.
Thus flux = Ky*(Yi-Yb)

Kx(Xb-Xi)=Ky*(Yi-Yb) equation 1
When we increase the amount of the gas without changing the liquid , bulk liquid concentration(Xb) will increase, Yi and Xi are in equilibrium, ( since Yi is not changing as we are not changing concentration in gas ) thus Xi will not change.
Since the left side of the equation will not change, we can say that Kx will decrease.

This changing the gas flow rate also changes the mass transfer cofficients in liquid phase.

Hope this helps

Thanks